There are 12,000 people at a BTS concert. The ratio of children to adults is 2 to 4. Exactly how many more adults are there than children at the concert? Show all work and label your answer


1
Expert's answer
2021-03-28T17:16:46-0400

Given that,

There are 12,000 people at a BTS concert. The ratio of children to adults is 2 to 4.

Let C be the number of children in the concert and A be the number of adults in the concert.

"\implies" given:

A + C = 12000 "\to" (1)

also given:

"\frac{C}{A}" = "\frac{2}{4}"


"\implies"2A = 4C

"\implies"2A - 4C = 0

"\implies"A - 2C = 0 "\to" (2)

by solving (1) and (2) we get:

A = 8000 & C = 4000

Hence, exactly 4000 more adults are there than children at the concert.